OpenStudy (anonymous):

My result was x=51 but the test answer is x=75. 5x-4/7=2x+9/3. which is correct

OpenStudy (anonymous):

(5x-4)/7=(2x+9)/3 multiply each side by 21 (21/7)(5x-4)=(2x+9)(21/3) 3(5x-4)=7(2x+9) 15x-12=14x+63 x=75 now check (5*75-4)/7=(2*75+9)/3 (375-4)/7=(150+9)/3 371/7=159/3 53=53 now check ur answer: (5*51-4)/7=(2*51+9)/3 (255-4)/7=(102+9)/3 251/7=111/3